OpenStudy (ttop0816):

help! Find the phase shift of the function y = 5cos(2x +pi/2 ). left pi/2 <<< i chose this! right pi/2 left pi/4 right pi/4

OpenStudy (anonymous):

it is \[ \frac \pi 4 \]

OpenStudy (ttop0816):

how come??

OpenStudy (perl):

factor the argument of your trig function

OpenStudy (anonymous):

cos( x+a) has phase a cos(2 x + a) has phase a/2

OpenStudy (phi):

match your equation to \[ A \cos\left( \omega(x + \phi)\right) \] to make your equation look like that, you must factor a 2 from (2x + pi/2) to do this write it as (2x + 2/2 * pi/2) now factor out 2: 2(x+½ pi/2) or 2(x+ pi/4)

OpenStudy (perl):